You add 200.0 mL of 0.10 M KOH to 50.0 mL of 0.40 M HF. What is the pH of the resulting solution? Provide answer with solution.


1
Expert's answer
2021-04-16T04:59:49-0400

KOH + HF = KF + H2O

CM = n / V

n = CM x V

According to the equation, n (KOH) = n (HF).

n (KOH) = 0.10 x 0.2000 = 0.02 mol

n (HF) = 0.40 x 0.0500 = 0.02 mol

In this case, the base is completely neutralized by acid. Therefore, the resulting pH=7.